OpenStudy (anonymous):

find the x and y intercept of the line -10x-6y=120

OpenStudy (anonymous):

Please help me!!!

OpenStudy (accessdenied):

x-intercept is the point(s) where the graph touches the x-axis The x-axis is the line y=0, so we can plug this in and solve for x. Our points will look like this: (x, 0) (x units left/right, 0 units up/down). y-intercept is the point(s) where the graph touches the y-axis. The y-axis is the line x=0, so we can plug this in and solve for y. Our points will look like this: (0, y) (0 units left/right, y units up).

OpenStudy (anonymous):

i tried to solve it, and got it wrong. I thought it was x-intercept=-20 and y-intercept=-12... where did i go wrong and what's the right answer??

OpenStudy (accessdenied):

Could you show some work with those results? That's be helpful to see where you got those answers from. :)

OpenStudy (anonymous):

-10x-6y=120 -10(0)-6y=120 -10*0=0 -6y=120 120/-6=-20

OpenStudy (anonymous):

i got -12 by -10x-6(0)=120 -6*0=0 -10x=120 120/-10=-12

OpenStudy (accessdenied):

So, that's the point (0, -20). That would be a y-intercept. And your second one is (-12, 0), which is an x-intercept.

OpenStudy (anonymous):

i switched it????!!!!

OpenStudy (accessdenied):

Yes, it appears so. :P y-intercepts are where you plug x=0 in: (0, y). x-intercepts are where you plug y=0 in: (x, 0). It can be confusing at times if you're not careful. :)

OpenStudy (anonymous):

thank you! how do i best answer you?

OpenStudy (anonymous):

nevermind

OpenStudy (accessdenied):

You're welcome! :)